OpenStudy (anonymous):

A spinner is divided into many sections of equal size. Some sections are red, some are blue, and the remaining are green. The probability of the arrow landing on a section colored red is 11 over 20. The probability of the arrow landing on a section colored blue is 6 over 20. What is the probability of the arrow landing on a green-colored section? 6 over 20 3 over 20 9 over 20 17 over 20 @thesmartone

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

P(red) = 11/20 P(blue) = 6/20 P(green) = x

jimthompson5910 (jim_thompson5910):

P(red) + P(blue) + P(green) = 1 11/20 + 6/20 + x = 1 solve for x

OpenStudy (anonymous):

jim im confused ?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is 11/20 + 6/20 equal to

OpenStudy (anonymous):

17/20 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so we have 17/20 + x = 1

jimthompson5910 (jim_thompson5910):

subtract 17/20 from both sides

jimthompson5910 (jim_thompson5910):

x = 1 - (17/20) x = 20/20 - 17/20 ... think of 1 as 20/20 x = ???

OpenStudy (anonymous):

so 1/1

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

3/20 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yep, so that's P(green)

OpenStudy (anonymous):

ok one last question